OpenStudy (anonymous):

How can you factor an algebraic expression in terms of a single trig function

hartnn (hartnn):

any particular example question u have ?

OpenStudy (anonymous):

sin^2 x+sin x-2

hartnn (hartnn):

yeah, so just assume sin x = a and u get an algebraic expression, factor that.

hartnn (hartnn):

then after factoring, you can put back a = sin x.

OpenStudy (anonymous):

my teacher keeps telling me I am overthinking it I got sinxsinx+sinx-2 /sinx+xinx =sinx-2

OpenStudy (anonymous):

what is a common factor

hartnn (hartnn):

a^2+a-2 = a^2 +2a-a-2 = a(a+2)-1(a+2) did u get that ?

OpenStudy (anonymous):

ah ok, brain fart.

hartnn (hartnn):

that comes out to be (a-1)(a+2) and putting back a=sin x u get (sin x-1)(sin x+2)

OpenStudy (anonymous):

ok, makes so much more sense now.

hartnn (hartnn):

yeah, thats the final factored form

OpenStudy (anonymous):

ok thank you

hartnn (hartnn):

welcome ^_^